OpenStudy (marmar10):

A solution is prepared that contains a mixture of acetic acid and sodium acetate. The pKa of acetic acid is 4.76 and the pH of the prepared solution is 4.21. What does this information tell you about the relative concentrations of acetic acid and sodium acetate?

OpenStudy (explorerhills):

this is a buffer solution of weak acid and its salt. so use the formula that we use to find the ph of a buffer solution.

OpenStudy (marmar10):

what is that formula called?

OpenStudy (cuanchi):

@NELearner in the equation the ratio is the other way the base over the acid (salt/acid) or you have to change the sign toa a negative sign

OpenStudy (anonymous):

@Cuanchi Good grief, my apologies, everyone.

OpenStudy (anonymous):

We are told that the pKa of acetic acid is 4.76. Remember, the smaller the value of pKa, the stronger the acid. The Henderson-Hasselbalch Equation, which explorerhills referenced, is as follows: \[pH=pKa+\log \left( \frac{ [CH₃COONa] }{ [CH₃COOH]} \right)\] Subtracting pKa from both sides: \[pH-pKa=\log \left( \frac{ [CH₃COONa] }{ [CH₃COOH]} \right)\] Since we know the pH and the pKa, we can substitute and get a ratio of acid to base: \[4.21-4.76 = \log \left( \frac{ [CH₃COONa] }{ [CH₃COOH]} \right)\] This ratio will give you information about the relative concentrations of acetic acid and sodium acetate.